OpenStudy (anonymous):

Simplify (21+3x/5x^4)/(49-x^2/7x)

OpenStudy (anonymous):

3(x5+35)5(49−x 97

OpenStudy (mathstudent55):

Do you mean a complex fraction? \(\Large \dfrac{ \frac{21 + 3x}{5x^4}}{\frac{49 - x^2}{7x}} \)

OpenStudy (anonymous):

Or go to a website called : www.mathway.com

OpenStudy (anonymous):

alright ill check it out and yeah a complex fraction just like that

OpenStudy (mathstudent55):

Remember, a complex fraction is a fraction divided by a fraction. To divide fractions, multiply the first one by the reciprocal of the second one. \(\Large \dfrac{ \frac{21 + 3x}{5x^4}}{\frac{49 - x^2}{7x}}\) \(= \Large { \frac{21 + 3x}{5x^4}} \div {\frac{49 - x^2}{7x}} \) \(= \Large { \frac{21 + 3x}{5x^4}} \times {\frac {7x} {49 - x^2} } \) Now factor every numerator and denominator and simplify.

OpenStudy (anonymous):

would this be my answer

OpenStudy (mathstudent55):

\( = \Large { \frac{21 + 3x}{5x^4}} \times {\frac {7x} {49 - x^2} }\) \( = \Large { \frac{3(7 + x)}{5x^4}} \times {\frac {7x} {(7 + x)(7 - x)} }\) \(= \Large { \frac{3\cancel{(7 + x)}}{5x^4}} \times {\frac {7x} {\cancel{(7 + x)}(7 - x)} }\) \(= \Large {\frac {21\cancel{x}} {5x^3\cancel{x}(7 - x)} }\) \(= \Large {\frac {21} {5x^3(7 - x)} }\)